OpenStudy (anonymous):

I don't get this, can someone help? "Write the equation of a circle with endpoints of the diameter at (4, -3) and (-2, 5)."

OpenStudy (zzr0ck3r):

so the distance of points \((x_0,y_0)\) and \((x_1,y_1)\) is given by \(\sqrt{(x_0-x_1)^2+(y_0-y_1)^2}\) this will also be the diameter. the equation of a circle at center \((a,b)\) is \((x-a)^2+(x-b)^2=r^2\) where \(r\) is the radius of the circle. So using the formula above we can get the diameter and then we have the radius, so we need only to find the center point \((a,b)\). you with me so far?

OpenStudy (zzr0ck3r):

@Bwaulioo ?

OpenStudy (zzr0ck3r):

that equation of a circle should say \((x-a)^2+(y-b)^2=r^2\)

OpenStudy (anonymous):

yeah sorry >.<

OpenStudy (zzr0ck3r):

the center will be given by \((\frac{4+(-2)}{2},\frac{-3+5}{2})\)

OpenStudy (zzr0ck3r):

so \(r=\frac{1}{2}\sqrt{(4-(-2))^2+(-3-5)^2}=\frac{1}{2}\sqrt{36+64}=\frac{1}{2}\sqrt{100}=5\) and our center is \((1,1)\) so our equation is \((x-1)^2+(y-1)^2=25\)

OpenStudy (anonymous):

ohhhhhh, I get it, wow I knew part of that but dind't know exact;y how to go about, Thanks so much!

OpenStudy (zzr0ck3r):

np